OpenStudy (anonymous):

Need help with the final step:: Let the function f(x) be defined by the equation y = f(x), where x and f(x) are real numbers. f(x) = 4 − x^2 f(9) = -77 f(-7) = -45 f(k) = 4-k^2 f(k2 − 8) = ??

OpenStudy (anonymous):

Basically you just plug in the entire input expression in place of the x so f(k2 - 8) = 4 - (k2 - 8)^2

OpenStudy (anonymous):

suppose to have read as f(k^2-8)=

OpenStudy (anonymous):

Still valid if you read the answer the same way. =) f(k^2 - 8) = 4 - (k^2 - 8)^2